OpenStudy (kyleighg):

i need help on factoring 2t^2+5t-3

OpenStudy (mathmale):

'Which methods of factoring quadratics have you used in the past? 1) completing the square 2) factoring by grouping 3) synthetic division 4) Rational Root Theorem Try one or more of these.

OpenStudy (kyleighg):

can you help me with the factoring by grouping?

OpenStudy (mathmale):

We could certainly try that. Unfortunately, factoring by grouping would be difficult to apply here, although we might be able to verify factors found in other ways.

OpenStudy (mathmale):

I just named several methods of factoring quadratics, not meaning to recommend any of them by the order i which I presented them.

OpenStudy (kyleighg):

i would just like to know how to do one of them please

OpenStudy (mathmale):

Try this: Take the first coefficient (2) and the last coefficient (-3) and multiply them together. What are factors of the result? (2)(-3)=?

OpenStudy (kyleighg):

-6

OpenStudy (kyleighg):

So now what do i do @mathmale

OpenStudy (mathmale):

List all of the possible factors of -6: \[\pm 6;~\pm 3;~\pm 2;~\pm 1\]

OpenStudy (mathmale):

If you know synthetic div, I'd strongly recommend that you use it to determine whether or not any of those 8 possible roots are indeed solutions of the given quadratic. If y ou don't know synth. div., there are other methods, such as long division.

OpenStudy (kyleighg):

i am quite lost

OpenStudy (mathmale):

All right. Let's revert to the quadratic formula (which I should have listed earlier). Heard of and/or used that?

OpenStudy (kyleighg):

My teacher was going to explain that later this week but i dont mind giving it a try

OpenStudy (mathmale):

\[If~ ax^2+bx+c=0,~x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (mathmale):

You are given the quadratic equation 2x^2 + 5x - 3 = 0, for which the coefficients a, b and c are 2, 5 and -3 respectively. Try putting those values into the formula for x and calculating x that way.

OpenStudy (kyleighg):

so would my equation be x=-5+/- the square root of 5^2-4 over 2(2)

OpenStudy (mathmale):

Good start. Please, could you put that into a nicer mathematical format (e. g., using a horiz line instead of "over")? Simplify 5^2 - 4.

OpenStudy (kyleighg):

Well i would if i could

OpenStudy (mathmale):

so would my equation be x=-5+/- the square root of 5^2-4 over 2(2)\[x=\frac{ -5\pm \sqrt{5^2-4} }{ 2(2) }\]

OpenStudy (mathmale):

Check your "b^2-4ac." I think you've forgotten your "a" and "c".

OpenStudy (mathmale):

b=5; a=2; and c=-3. What is the "determinant," b^2 - 4ac? ( )^2 - 4 ( )( )

OpenStudy (kyleighg):

so is this what it should look like

OpenStudy (mathmale):

Were you kicked off the 'Net or did you log off? If you log off in mid-stream, please let me know your intentions beforehand. I really wanted to help you.

OpenStudy (mathmale):

Very good. What is 5^2? What is 4(2)(-3)?

OpenStudy (mathmale):

OpenStudy first told me you were "offline" and then that you were "just looking around."

OpenStudy (mathmale):

What is the value of 5^2 - 4(2)(-3)?

OpenStudy (kyleighg):

i was online and i got the app

OpenStudy (kyleighg):

I got 1

OpenStudy (mathmale):

5^2 = 25 -4 (2) (-3) = ??

OpenStudy (mathmale):

Yes, 25-24 =1.

OpenStudy (mathmale):

So your two roots are:\[x=\frac{ -5\pm \sqrt{1} }{ 4 }\]

OpenStudy (mathmale):

sorry